# Magneto: Combining Small and Large Language Models for Schema Matching

### v1

@@ 7. Related Work
A wide range of methods have been proposed for schema matching, an important research area in the database field (Rahm and Bernstein, 2001). Related approaches to our work can be broadly categorized into three main groups: traditional methods, small language model-driven methods, and large language model-based methods. Below, we review representative work in each category.
@@
Traditional Methods. Over the years, researchers have developed and combined various approaches for automatic schema matching (Bellahsene et al., 2011; Doan et al., 2012; Koutras et al., 2021). A straightforward yet effective approach is detecting overlap in column names (Madhavan et al., 2001; Melnik et al., 2002; Do and Rahm, 2002) and overlap in column values (Aumueller et al., 2005; Zhang et al., 2011). Some incorporate relaxations when measuring overlaps, such as accounting for syntactic and semantic similarities between column names and values (Madhavan et al., 2001; Aumueller et al., 2005). Others also consider factors like data type relevance and value distribution (Aumueller et al., 2005; Zhang et al., 2011). Among these, the COMA algorithm stands out for integrating the most strategies and weighting their outputs to achieve better accuracy (Do and Rahm, 2002; Aumueller et al., 2005), often remaining competitive even against more recent approaches (Koutras et al., 2021).
However, these approaches often struggle to capture complex relationships and deeper semantics within datasets, resulting in poor accuracy when identifying semantically valid matches (Khatiwada et al., 2023).
Small Language Model-Driven Methods. Small Language Models (SLMs) have transformed schema matching by using embeddings similarity (Cappuzzo et al., 2020), effectively encoding and comparing textual and contextual data within columns. Furthermore, contrastive learning refines SLMs’ ability to distinguish between matching and non-matching column pairs (Fan et al., 2023; Cong et al., 2023; Zhang et al., 2023). Meanwhile, synthetic tabular data generation facilitates self-supervised learning, allowing models to improve without ground truth (Fan et al., 2023; Du et al., 2024).
@@
Among these methods, ISResMat customizes pre-trained models for dataset-specific adaptation, generating training pairs from table fragments, and applying pairwise comparison losses to refine matching accuracy (Du et al., 2024). It also addresses cardinality constraints and manages numerical columns with advanced techniques; Unicorn, a state-of-the-art general matching model, employs contrastive learning and a Mixture-of-Experts (MoE) layer within its architecture to discern matches but relies on supervised training (Tu et al., 2023).
@@
Large Language Model-Driven Methods. Recent works have leveraged large language models (LLMs) for various aspects of tabular data management, predominantly focusing on single-table tasks (Kayali et al., 2024; Li et al., 2024a; Feuer et al., 2024; Hegselmann et al., 2023). Some studies discuss LLM applications in schema matching, highlighting the potential of LLMs in this area. However, these approaches often rely solely on prompting strategies—either fine-tuned or zero-shot—which suffer from scalability issues and high computational costs (Li et al., 2024a; Parciak et al., 2024; Fang et al., 2024; Feng et al., 2024).
Furthermore, Sheetrit et al. (2024) and Zhang et al. (2024) utilize zero-shot pre-trained LLMs for schema matching. Sheetrit et al. (2024) addresses multiple-table matching, whereas our work focuses on the less-explored two-table schema matching, particularly for tables with numerous columns. Zhang et al. (2024) uses rule-based feature extraction and trains an XGBoost classifier with gold data, a supervised approach distinct from our unsupervised method. Hence, we do not compare our findings directly with these studies.
8. Conclusions
@@
We have proposed Magneto, a novel framework that leverages small and large language models to derive schema matching strategies that generalize across domains and balance accuracy and runtime tradeoffs. We introduce a new benchmark that captures some of the complexities involved in biomedical data integration and that also presents new challenges for schema matching methods. Through a detailed experimental evaluation, including comparisons against state of the art methods and ablations, we demonstrate the effectiveness of Magneto and our design choices.
The source code for all baselines and Magneto can be accessed from https://github.com/VIDA-NYU/data-integration-eval and https://github.com/VIDA-NYU/data-harmonization-benchmark.
Acknowledgments.
@@
We thank our collaborators from the NYU School of Medicine, David Fenyo, Wenke Liu, and Sarah Keegan, for sharing their experiences in biomedical data integration and contributing to the creation of the GDC benchmark. We would like to thank DARPA for funding this work through the Automating Scientific Knowledge Extraction and Modeling (ASKEM) program, Agreement No. HR0011262087. This work was also partially supported by NSF award IIS-2106888. The views, opinions, and findings expressed are those of the authors and should not be interpreted as representing the official views or policies of the Department of Defense, the U.S. Government, or NSF. We also thank Roque Lopez, Benjamin Feuer, and Grace Fan for their constructive suggestions.

### v2

@@ 7. Related Work
This section discusses schema and ontology matching approaches related to our work.
@@
Traditional Methods. A straightforward approach for schema matching is detecting overlap in column names (Madhavan et al., 2001; Melnik et al., 2002; Do and Rahm, 2002) and overlap in column values (Aumueller et al., 2005; Zhang et al., 2011). Some incorporate relaxations when measuring overlaps, such as accounting for syntactic and semantic similarities between column names and values (Madhavan et al., 2001; Aumueller et al., 2005). Others also consider factors like data type relevance and value distribution (Aumueller et al., 2005; Zhang et al., 2011). Among these, the COMA algorithm stands out for integrating the most strategies and weighting their outputs to achieve better accuracy (Do and Rahm, 2002; Aumueller et al., 2005), often remaining competitive even against more recent approaches (Koutras et al., 2021). However, these approaches often struggle to capture complex relationships and deeper semantics within datasets (Khatiwada et al., 2023).
@@
Small Language Model-Driven Methods. Methods based on small language models (SLMs) usually use embeddings to encode and compare column data (Cappuzzo et al., 2020). Contrastive learning can improve an SLM’s ability to distinguish matching and non-matching column pairs (Fan et al., 2023; Cong et al., 2023; Zhang et al., 2023), while synthetic tabular data generation can help models to improve without ground truth (Fan et al., 2023; Du et al., 2024). Among these methods, ISResMat customizes pre-trained models for dataset-specific adaptation, generating training pairs from table fragments, and applying pairwise comparison losses to refine matching accuracy (Du et al., 2024). Unicorn, a general matching model, employs contrastive learning and a Mixture-of-Experts (MoE) layer within its architecture to discern matches but relies on supervised training (Tu et al., 2023).
@@
Large Language Model-Driven Methods. Recent works have leveraged large language models (LLMs) for various aspects of tabular data management, predominantly focusing on single-table tasks (Kayali et al., 2024; Hegselmann et al., 2023). Some studies discuss LLM applications for schema matching and highlight the potential of LLMs for this task (Li et al., 2024a; Feuer et al., 2024). However, these approaches often rely solely on prompting strategies—either fine-tuned or zero-shot—which suffer from scalability issues and high computational costs (Li et al., 2024a; Parciak et al., 2024; Fang et al., 2024; Feng et al., 2024).  Sheetrit et al. (2024) and Zhang et al. (2024) utilize zero-shot pre-trained LLMs for schema matching. Sheetrit et al. (2024) addresses multiple-table matching, whereas our work focuses on two-table schema matching, particularly for tables with numerous columns. Zhang et al. (2024) uses rule-based feature extraction and trains an XGBoost classifier with gold data, a supervised approach distinct from our unsupervised method. Xu et al. (2024) proposed manually-derived rules to guide LLMs during the matching process and incorporated external knowledge to deal with hallucinations. Parciak et al. (2024) explored various prompting strategies for matching source attributes to a target schema. Note that the evaluation of these methods only take into account table/attribute names and descriptions (Xu et al., 2024; Parciak et al., 2024)—unlike Magneto, they do not consider values.
Ontology Matching. Ontology matching (OM) is related to schema matching (SM) but focuses on identifying semantic correspondences between elements (e.g., classes and properties) across different ontologies expressed in languages like OWL and RDF/XML (Euzenat and Shvaiko, 2013). Some OM approaches use heuristics, rule-based methods (Euzenat and Shvaiko, 2013; Hertling et al., 2019), structural similarity, linguistics, and domain-specific resources (Jiménez-Ruiz et al., 2012; Harrow et al., 2017). To facilitate the matching of domain-specific terminology, semi-automatic annotation using vocabularies and ontologies has been used to enrich schema labels (Sorrentino et al., 2011; Beneventano et al., 2015). Recent OM approaches leverage LLMs and, like Magneto, use a two-phase matching process (Babaei Giglou et al., 2025; Hertling and Paulheim, 2023). OLaLa uses embeddings for candidate matches and LLM evaluation with natural language conversion (Hertling and Paulheim, 2023), while LLMs4OM employs RAG to extract and classify concept similarities (Babaei Giglou et al., 2025). Unlike these methods that prompt per candidate pair, Magneto efficiently generates a ranked list for each input column.
Adapting OM methods to SM is not straightforward: OM approaches target structured relationships rather than flat tabular data, may disregard the similarity of column values, and require format conversions that can impact performance. We compared Magneto against two OM systems, LogMap (Jiménez-Ruiz et al., 2012) and LLMs4OM (Babaei Giglou et al., 2025), by treating tables as classes and columns as properties. They underperformed SM-based baselines by a large margin. Nonetheless, given their shared goals and challenges, combining OM and SM techniques is an interesting direction for future work.
8. Conclusions and Future work
@@
We proposed Magneto, a framework that leverages small and large language models to derive schema matching strategies that generalize across domains and balance accuracy and runtime tradeoffs. We introduced a new benchmark that captures some of the complexities in biomedical data integration and presents new challenges for schema matching. With a detailed experimental evaluation, including comparisons against state-of-the-art methods and ablations, we demonstrate the effectiveness of Magneto and our design choices.
There are several directions for our future work. Magneto ’s accuracy depends on SLM retrieval—if the correct match is missing from the top-$k$, reranking cannot recover it. We plan to explore hybrid retrieval strategies and training SLMs on large, multi-domain datasets. To address limitations in underrepresented domains, we also aim to integrate external knowledge for better zero-shot performance. Reranking quality can vary due to prompt dependence—a known limitation of instruction-tuned LLMs (Feuer et al., 2024). Future work may leverage prompt tuning frameworks like DSPy (Khattab et al., 2024) for systematic, task-aware optimization. While LLM-based reranking improves accuracy, it introduces cost overheads. Future work could optimize reranking with smaller, efficient LLMs.
Acknowledgments
@@
This work was supported by NSF awards IIS-2106888 and OAC-2411221, the DARPA ASKEM program Agreement No. HR0011262087, and the ARPA-H BDF program. The views, opinions, and findings expressed are those of the authors and should not be interpreted as representing the official views or policies of the DARPA, ARPA-H, the U.S. Government, or NSF. We thank our collaborators from the NYU School of Medicine, David Fenyo, Wenke Liu, and Sarah Keegan, for sharing their experience in biomedical data integration and contributing to the GDC benchmark. We also thank Roque Lopez for his contributions to the code.